OpenStudy (anonymous):

i will give out medals @Here_to_Help15

OpenStudy (anonymous):

here

OpenStudy (anonymous):

@jagr2713 do u know it

OpenStudy (anonymous):

@Here_to_Help15 @ilovewolf

OpenStudy (anonymous):

@WalkerSlayer @shinebrightlikeadimon @gelow @jordanloveangel @xo.minnie.xox

OpenStudy (anonymous):

please help guys

OpenStudy (jordanloveangel):

ur doing study island right ?

OpenStudy (anonymous):

yes

OpenStudy (jordanloveangel):

k m i try

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@King.Void.

OpenStudy (jordanloveangel):

np

OpenStudy (anonymous):

It would be B- 22-6=16 16*x=16x 1*8=8

OpenStudy (anonymous):

no not B

OpenStudy (jordanloveangel):

8(2x + 1)

OpenStudy (anonymous):

yes can u do some more please

OpenStudy (anonymous):

here

OpenStudy (jordanloveangel):

im not good at those

OpenStudy (anonymous):

@K12_math

OpenStudy (anonymous):

@AndrewTheCookie

OpenStudy (anonymous):

do you still need help?

OpenStudy (anonymous):

4x+8/3

OpenStudy (anonymous):

can u do this one

OpenStudy (anonymous):

37x-108